OpenStudy (anonymous):

Let f(x) = \sqrt{20 - x} and g(x) = x^2 - x. Then the domain of f\circ g is equal to [a,b] for

OpenStudy (anonymous):

\[(f\circ g)(x) = \sqrt{20-\left( x^2 - x \right)}\Rightarrow \left[ 20-\left( x^2 - x \right) \right] \ge 0\]

OpenStudy (anonymous):

I need in [a,b], infinity is not defined

OpenStudy (anonymous):

you just looking for an answer?

OpenStudy (anonymous):

yeah I am just looking for an answer because the answers I keep getting are wrong.

OpenStudy (anonymous):

can you solve the last inequality I wrote?

OpenStudy (anonymous):

Yes I got it now THANK YOU!

OpenStudy (anonymous):

you're welocme!

OpenStudy (anonymous):

welcome

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.